OpenStudy (anonymous):

This is for a 30 60 90 triangle that has only given me the long leg which is 3 i need to find the hypotnuse. please help

OpenStudy (anonymous):

Can you create a diagram or upload a picture of the triangle to clarify?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it just says what i wrote

OpenStudy (anonymous):

it does give the third number which is 3

OpenStudy (anonymous):

I'll have to infer what the picture might look like then:  We can use the law of sines, which states that each side over the sin of their angle is constant throughout the triangle, so: \[\frac{ 3 }{ \sin 60 } = \frac{ n }{ \sin 90} so n=3(\sin 90)/(\sin 60) \approx 3.464\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the hypotenuse would be 3.464

jhonyy9 (jhonyy9):

do you have any choises ?

OpenStudy (anonymous):

no choices it just gives me the third number which is 3

jhonyy9 (jhonyy9):

so what mean the third number is 3 ?

OpenStudy (anonymous):

Yes, I solved for the hypotenuse, which is n.

OpenStudy (anonymous):

its a drag and drop online question for my class and well it just gives me the number 3

OpenStudy (anonymous):